# Features of the Partition Function of a $\Lambda>0$ Universe

### v1

@@ 3.2 Hodge-De Rham operator spectrum
The eight $\lambda^{(1)}_{1}$ eigenmodes are the Killing vector fields of the Fubini-Study metric. We note that $\Delta_{(1)}=-\nabla^{2}_{(1)}+\Lambda$, where ($\nabla^{2}_{(1)}$ is the Laplacian acting on transverse vector fields. The operator $\Delta_{(1)}$ acts on the transverse part of the gauge field in Maxwell theory on an Einstein metric background satisfying $R_{\mu\nu}=\Lambda g_{\mu\nu}$. The eigenmodes are associated with the $(n+1,n+1)\oplus(n-1,n+2)\oplus(n+2,n-1)$ irreps of $SU(3)$.
@@ 3.3 Lichnerowicz operator spectrum
We note that $\Delta_{(2)}h_{\mu\nu}=-\nabla^{2}_{(2)}h_{\mu\nu}-2R_{\mu\nu\alpha\beta}h^{\alpha\beta}+2\Lambda h_{\mu\nu}$, where ($\nabla^{2}_{(2)}$ is the Laplacian acting on transverse-traceless symmetric rank two tensors. The shifted operator $\Delta_{(2)}-2\Lambda$ appears in the linearised Einstein theory with $\Lambda>0$, expanded around an Einstein metric satisfying $R_{\mu\nu}=\Lambda g_{\mu\nu}$. The eigenmodes are associated with the $(n+2,n+2)\oplus(n+1,n+4)\oplus(n+4,n+1)\oplus(n,n+6)\oplus(n+6,n)$ irreps of $SU(3)$. [^7] We have corrected a minor typographical error in the left most column of Table 1 of Warner:1982fs .
@@ 4.1 Heat kernel method
[…]
@@
[…]
@@ 4.3 Comparison to $S^{4}$
Here we have defined $\tilde{\nu}\equiv\sqrt{3\mu^{2}\Lambda^{-1}-\tfrac{9}{4}}$. Upon formally setting $\varepsilon=0$, the integrand of above expression contains the $SO(4,1)$ Harish-Chandra character for the principal series irreducible representation labelled by the quadratic Casimir $\mathcal{C}_{\tilde{\nu}}\equiv\Delta_{\tilde{\nu}}(3-\Delta_{\tilde{\nu}})$ with $\Delta_{\tilde{\nu}}=\tfrac{3}{2}+i\tilde{\nu}$. The Harish-Chandra character reads
@@
Here $\rho_{\tilde{\nu}}(\omega)$ is a density of states obtained by Fourier transforming $\chi_{\Delta}(t)$. When evaluating the heat kernel integral (4.18), Hurwitz zeta functions of the type $\zeta^{\prime}(-n,\tfrac{3}{2}\pm i\tilde{\nu})$ with $n=0,1,2,3$ appear (as an example, see (C.27) with $g_{s}=1$ in Anninos:2020hfj ).
@@ 4.4 Non-perturbative features
$\mathcal{Z}_{\text{grav}}\approx e^{\mathcal{S}_{S^{4}}}\mathcal{Z}^{(1)}_{S^{4}}\mathcal{Z}^{(\varphi)}_{S^{4}}+e^{\mathcal{S}_{\mathbb{C}P^{2}}}\mathcal{Z}^{(1)}_{\mathbb{C}P^{2}}\mathcal{Z}^{(\varphi)}_{\mathbb{C}P^{2}}+\ldots$ (4.23)
@@
We can view the contribution from $\mathbb{C}P^{2}$ as a non-perturbatively small correction to the leading $S^{4}$ contribution. The contribution from $\mathbb{C}P^{2}$ cannot be expressed as a $SO(4,1)$ Harish-Chandra character, so we interpret it as a small de Sitter symmetry breaking effect of size $e^{\mathcal{S}_{\mathbb{C}P^{2}}-\mathcal{S}_{S^{4}}}=e^{-\tfrac{3\pi}{4\Lambda G}}$. For the current value of the cosmological constant, this is an extraordinarily small number of the order $e^{-10^{122}}$. [^8] In particular, it is exponentially smaller than the putative effects discussed in Tsamis:1996qq ; Polyakov:2007mm . It would be interesting to understand if it could compete with other non-perturbative instabilities, for instance when de Sitter space is realised as a metastable vacuum in string theory. If we also add the contribution from higher-derivative terms, such as the Gauss-Bonnet term or the Weyl squared term, the suppression factor will also depend on the value of their respective couplings. We expect higher-derivative effects to be suppressed by some high energy scale in our present cosmology, but perhaps during primordial inflation they could give rise to more enhanced contributions.
@@ 5.4 Comparing to the heat kernel coefficient
We note that, as detailed in appendix G of Anninos:2020hfj , the final $-2$ stems from a proper treatment of the path integration measure. It is precisely the $\Lambda_{\text{u.v.}}^{-1}\sim\varepsilon^{2}$ prefactor appearing in (5.6). A similar treatment can be employed for the Maxwell theory on a round $S^{4}$, see for example Giombi:2015haa ; Anninos:2020hfj ; RiosFukelman:2023mgq , which yields a coefficient
@@ 6.1 Linearised action and measure
We note the kinetic term in the action $S_{E}^{(T)}$ has the wrong sign. This is the linearised version of the conformal mode problem, and we will treat it along the lines of Gibbons:1978ac . Namely we will rotate the contour $\tilde{h}\to i\tilde{h}$ and rotate back any residual negative modes. [^9] It was recently argued Ivo:2025yek that a more sophisticated continuation prescription for the overall phase is required to ensure gauge invariance of the path integral. We will not keep track of the overall sign of the phase to this degree of precision throughout this section. Due to the absence of non-isometric conformal Killing vectors, the only negative mode is in fact the constant mode of $\tilde{h}$. This also follows from a theorem of Lichnerowicz and Obata lichnerowicz1958geometrie ; obata1971conjectures . Thus, the phase of the one-loop partition function is ±$\pm i$.
@@ 6.4 Heat kernel integrals
~$\tilde{\mathcal{I}}_{11}=\int_{\varepsilon}^{\infty}\frac{dt}{\sqrt{t^{2}-\varepsilon^{2}}}\frac{-8e^{-2t}+31e^{-t}+27e^{t}-44}{\left(e^{t}-1\right)^{4}}\times\cosh\left(2\sqrt{t^{2}-\varepsilon^{2}}\right)~{}.$ (6.25)
@@
~$\tilde{\mathcal{B}}_{11}=-\frac{599}{120}~{},$ (6.26)
associated with the logarithmic divergence. The heat kernel integral ~$\tilde{\mathcal{I}}_{11}$ is evaluated in appendix B.2, and the resulting expression is provided in (B.4).
@@
~$\tilde{\mathcal{I}}_{12}=\int_{\varepsilon}^{\infty}\frac{dt}{\sqrt{t^{2}-\varepsilon^{2}}}\frac{2e^{\frac{t}{2}}(9\sinh t+11\cosh t-5)}{\left(e^{t}-1\right)^{4}}\times\cosh\left(\tfrac{\sqrt{13}}{2}\sqrt{t^{2}-\varepsilon^{2}}\right)~{}.$ (6.27)
@@
~$\tilde{\mathcal{B}}_{12}=-\frac{7}{30}~{},$ (6.28)
associated to the logarithmic divergence. The heat kernel integral ~$\tilde{\mathcal{I}}_{12}$ is evaluated in appendix B.2, and the resulting expression is provided in (B.5).
@@
Once again, the relevant eigenvalues are a shifted version of (3.6), (3.7), and (3.8). Specifically, we must shift the eigenvalues by $-2\Lambda$. Even after the shift, all eigenvalues remain positive, indicating that the tensor fluctuations of $\mathbb{C}P^{2}$ are stable. This is similar to the tensor spectrum of $S^{4}$, but different from that of $S^{2}\times S^{2}$ and the Page metric, both of which exhibit one negative mode Volkov:2000ih ; Hennigar:2024gbg .
@@
~$\tilde{\mathcal{I}}_{21}=\int_{\varepsilon}^{\infty}\frac{dt}{\sqrt{t^{2}-\varepsilon^{2}}}\frac{-8e^{-2t}+31e^{-t}+27e^{t}-44}{\left(e^{t}-1\right)^{4}}\times\cosh\left(2\sqrt{t^{2}-\varepsilon^{2}}\right)~{}.$ (6.32)
@@
~$\tilde{\mathcal{B}}_{21}=-\frac{599}{120}~{}.$ (6.33)
The heat kernel integral ~$\tilde{\mathcal{I}}_{21}$ is evaluated in appendix B.2, and the resulting expression is provided in (B.4).
@@
~$\tilde{\mathcal{I}}_{22}=\int_{\varepsilon}^{\infty}\frac{dt}{\sqrt{t^{2}-\varepsilon^{2}}}\frac{-20e^{-\frac{5t}{2}}+80e^{-\frac{3t}{2}}-118e^{-\frac{t}{2}}+70e^{\frac{t}{2}}}{\left(e^{t}-1\right)^{4}}\times\cosh\left(\tfrac{\sqrt{13}}{2}\sqrt{t^{2}-\varepsilon^{2}}\right)~{}.$ (6.34)
@@
~$\tilde{\mathcal{B}}_{22}=-\frac{607}{30}~{}.$ (6.35)
The heat kernel integral ~$\tilde{\mathcal{I}}_{22}$ is evaluated in appendix B.2, and the resulting expression is provided in (B.6).
@@
~$\tilde{\mathcal{I}}_{23}=\int_{\varepsilon}^{\infty}\frac{dt}{\sqrt{t^{2}-\varepsilon^{2}}}\frac{4e^{-t}\left(9\sinh t+19\cosh t-16\right)}{\left(e^{t}-1\right)^{4}}\times\cosh\left(\sqrt{t^{2}-\varepsilon^{2}}\right)~{}.$ (6.36)
@@
~$\tilde{\mathcal{B}}_{23}=\frac{1741}{60}~{}.$ (6.37)
The heat kernel integral ~$\tilde{\mathcal{I}}_{23}$ is evaluated in appendix B.2, and the resulting expression is provided in (B.8).
@@ 6.5 Comparing to the heat kernel coefficient
We note that, as is indicated in the $\Lambda_{\text{u.v.}}$ dependent pre-factor of (6.21), the final $-2\times 8+1$ stems from a proper treatment of the path integration measure. A similar treatment can be done for the linearised Einstein theory about the round $S^{4}$ saddle, see for example Volkov:2000ih ; Anninos:2020hfj , which yields a coefficient
@@
where now $\text{vol}_{\text{can}}=\tfrac{2}{3}(2\pi)^{6}$ is the canonically normalised group volume of the $SO(5)$ isometry group of $S^{4}$. The overall phase of $\mathcal{Z}^{(1)}_{S^{4}}$ is (±$(\pm i)^{1+5}=-1$ Polchinski:1988ua .
@@ Acknowledgements
It is a great pleasure to acknowledge discussions with Damián Galante, Pietro Benetti Genolini, Thomas Hertog, Chris Herzog, Albert Law, Juan Maldacena, Miguel Montero, Beatrix Mühlmann, and Edgar Shaghoulian. D.A. is funded by the Royal Society under the grant “Concrete Calculables in Quantum de Sitter” and the STFC Consolidated grant ST/X000753/1. C.B. is funded by STFC under the grant reference STFC/2887726. S.B. is funded by STFC under the grant reference STFC/2928477.
@@ A.1 Comparison to coincident point limit of Green’s function
$-\frac{1}{2}\times V_{\mathbb{C}P^{2}}\times\lim_{\sigma\to 1}\partial_{\mu^{2}}^{(3)}G_{\mathbb{C}P^{2}}(\sigma)|_{\mu^{2}=1}=\partial_{\mu^{2}}^{(4)}\log\mathcal{Z}^{(\varphi)}_{\mathbb{C}P^{2}}|_{\mu^{2}=1}~{}.$ (A.9)
@@ A.2 Integral representation of Green’s function
In the above, ${X}$ and ${Y}$ are unit normal vectors on $\mathbb{R}^{6}$, $\sigma\equiv{X}\cdot{Y}$, and now we have
@@
$-\frac{1}{2}\times V_{\mathbb{C}P^{2}}\times\lim_{\sigma\to 1}\partial_{\mu^{2}}^{(3)}G_{\mathbb{C}P^{2}}(\sigma)|_{\mu^{2}=4}=\frac{3\zeta(5)}{256}~{},$ (A.15)
@@
It is interesting that the expression (A.14) takes a character-like form resembling the expressions in Anninos:2020hfj for the sphere. To do so it is convenient to use that $-\lambda\partial_{\lambda}\partial_{\nu}\tfrac{\lambda^{\pm i\nu}}{\log\lambda}=\nu\lambda^{\pm i\nu}$. Denoting ${\boldsymbol{\Delta}}=2+2i\nu$, with $\nu=\sqrt{\tfrac{1}{4}\mu^{2}-1}$, and upon integrating by parts in $\lambda$, we can re-express (A.14) as
$-\frac{1}{2}\times V_{\mathbb{C}P^{2}}\times G_{\mathbb{C}P^{2}}(\zeta^{i},\zeta^{i})=\frac{1}{8\nu}\partial_{\nu}\int_{0}^{1}\frac{d\lambda}{2\lambda(-\log\lambda)}\frac{1+4\lambda+\lambda^{2}}{1-\lambda}\frac{\lambda^{\frac{\boldsymbol{\Delta}}{2}}+\lambda^{\frac{\bar{\boldsymbol{\Delta}}}{2}}}{(1-\lambda)^{3}}~{}.$ (A.16)

### v3

@@ 3.2 Hodge-De Rham operator spectrum
The eight $\lambda^{(1)}_{1}$ eigenmodes are the Killing vector fields of the Fubini-Study metric. We note that $\Delta_{(1)}=-\nabla^{2}_{(1)}+\Lambda$, where $\nabla^{2}_{(1)}$ is the Laplacian acting on transverse vector fields. The operator $\Delta_{(1)}$ acts on the transverse part of the gauge field in Maxwell theory on an Einstein metric background satisfying $R_{\mu\nu}=\Lambda g_{\mu\nu}$. The eigenmodes are associated with the $(n+1,n+1)\oplus(n-1,n+2)\oplus(n+2,n-1)$ irreps of $SU(3)$.
@@ 3.3 Lichnerowicz operator spectrum
We note that $\Delta_{(2)}h_{\mu\nu}=-\nabla^{2}_{(2)}h_{\mu\nu}-2R_{\mu\nu\alpha\beta}h^{\alpha\beta}+2\Lambda h_{\mu\nu}$, where $\nabla^{2}_{(2)}$ is the Laplacian acting on transverse-traceless symmetric rank two tensors. The shifted operator $\Delta_{(2)}-2\Lambda$ appears in the linearised Einstein theory with $\Lambda>0$, expanded around an Einstein metric satisfying $R_{\mu\nu}=\Lambda g_{\mu\nu}$. The eigenmodes are associated with the $(n+2,n+2)\oplus(n+1,n+4)\oplus(n+4,n+1)\oplus(n,n+6)\oplus(n+6,n)$ irreps of SU(3)SU(3). [^8] We have corrected a minor typographical error in the left most column of Table 1 of Warner:1982fs .
@@ 4.1 Heat kernel method
$e^{-\left(\lambda_{n}^{(0)}+\mu^{2}\right)\tau}=\frac{e^{-(\mu^{2}-1)\tau}}{\sqrt{4\pi\tau}}\int_{\mathbb{R}}du\,{e^{-\frac{u^{2}}{4\tau}}}\,e^{i(n+1)u}~.$ (4.5)
@@
$\log\mathcal{Z}^{(\varphi)}_{\mathbb{C}P^{2}}=\int_{\mathbb{R}+i\delta}du\frac{e^{-\nu\sqrt{u^{2}+\varepsilon^{2}}}}{2\sqrt{u^{2}+\varepsilon^{2}}}\frac{e^{iu}\left(1+4e^{iu}+e^{2iu}\right)}{\left(1-e^{iu}\right)^{4}}~,$ (4.8)
@@ 4.3 Comparison to $S^{4}$
[…]
@@
[…]
@@ 4.4 Non-perturbative features
$\mathcal{Z}_{\text{grav}}\approx e^{\mathcal{S}_{S^{4}}}\mathcal{Z}^{(1)}_{S^{4}}\mathcal{Z}^{(\varphi)}_{S^{4}}+e^{\mathcal{S}_{\mathbb{C}P^{2}}}\mathcal{Z}^{(1)}_{\mathbb{C}P^{2}}\mathcal{Z}^{(\varphi)}_{\mathbb{C}P^{2}}$$\ldots$ (4.23)
@@
We can view the contribution from $\mathbb{C}P^{2}$ as a non-perturbatively small correction to the leading $S^{4}$ contribution. The contribution from $\mathbb{C}P^{2}$ cannot be expressed as a $SO(4,1)$ Harish-Chandra character, so we interpret it as a small de Sitter symmetry breaking effect of size $e^{\mathcal{S}_{\mathbb{C}P^{2}}-\mathcal{S}_{S^{4}}}=e^{-\tfrac{3\pi}{4\Lambda G}}$. For the current value of the cosmological constant, this is an extraordinarily small number of the order $e^{-10^{122}}$. [^9] In particular, it is exponentially smaller than the putative effects discussed in Tsamis:1996qq ; Polyakov:2007mm . It would be interesting to understand if it could compete with other non-perturbative instabilities, for instance when de Sitter space is realised as a metastable vacuum in string theory. If we also add the contribution from higher-derivative terms, such as the Gauss-Bonnet term or the Weyl squared term, the suppression factor will also depend on the value of their respective couplings. We expect higher-derivative effects to be suppressed by some high energy scale in our present cosmology. Perhaps during primordial inflation non-spherical topologies give rise to more enhanced contributions, particularly if they contribute to Lorentzian processes at sufficiently late-times. It is interesting to note, further, that in higher dimensions there exist saddles that break the de Sitter isometries boyer2005einstein ; bohm1998inhomogeneous on the topological sphere itself.
@@ 5.4 Comparing to the heat kernel coefficient
We note that, as detailed in appendix G of Anninos:2020hfj , the final 2-2 stems from a proper treatment of the path integration measure. It is precisely the $\Lambda_{\text{u.v.}}^{-1}\sim\varepsilon^{2}$ prefactor appearing in (5.6). A similar treatment can be employed for the Maxwell theory on a round $S^{4}$, see for example Giombi:2015haa ; Anninos:2020hfj ; RiosFukelman:2023mgq , which yields a coefficient
@@ 6.1 Linearised action and measure
We note the kinetic term in the action $S_{E}^{(T)}$ has the wrong sign. This is the linearised version of the conformal mode problem, and we will treat it along the lines of Gibbons:1978ac . Namely we will rotate the contour $\tilde{h}\to i\tilde{h}$ and rotate back any residual negative modes. [^10] It was recently argued Ivo:2025yek that a more sophisticated continuation prescription for the overall phase is required to ensure gauge invariance of the path integral. We will not keep track of the overall sign of the phase to this degree of precision throughout this section. Due to the absence of non-isometric conformal Killing vectors, the only negative mode is in fact the constant mode of $\tilde{h}$. This also follows from a theorem of Lichnerowicz and Obata lichnerowicz1958geometrie ; obata1971conjectures . Thus, the phase of the one-loop partition function is $\pm i$.
@@ 6.4 Heat kernel integrals
$\tilde{\mathcal{I}}_{11}=\int_{\varepsilon}^{\infty}\frac{dt}{\sqrt{t^{2}-\varepsilon^{2}}}\frac{-8e^{-2t}+31e^{-t}+27e^{t}-44}{\left(e^{t}-1\right)^{4}}\times\cosh\left(2\sqrt{t^{2}-\varepsilon^{2}}\right)~.$ (6.25)
@@
$\tilde{\mathcal{B}}_{11}=-\frac{599}{120}~,$ (6.26)
associated with the logarithmic divergence. The heat kernel integral $\tilde{\mathcal{I}}_{11}$ is evaluated in appendix B.2, and the resulting expression is provided in (B.4).
@@
$\tilde{\mathcal{I}}_{12}=\int_{\varepsilon}^{\infty}\frac{dt}{\sqrt{t^{2}-\varepsilon^{2}}}\frac{2e^{\frac{t}{2}}(9\sinh t+11\cosh t-5)}{\left(e^{t}-1\right)^{4}}\times\cosh\left(\tfrac{\sqrt{13}}{2}\sqrt{t^{2}-\varepsilon^{2}}\right)~.$ (6.27)
@@
$\tilde{\mathcal{B}}_{12}=-\frac{7}{30}~,$ (6.28)
associated to the logarithmic divergence. The heat kernel integral $\tilde{\mathcal{I}}_{12}$ is evaluated in appendix B.2, and the resulting expression is provided in (B.5).
@@
Once again, the relevant eigenvalues are a shifted version of (3.6), (3.7), and (3.8). Specifically, we must shift the eigenvalues by $2\Lambda$. Even after the shift, all eigenvalues remain positive, indicating that the tensor fluctuations of $\mathbb{C}P^{2}$ are stable. This is similar to the tensor spectrum of $S^{4}$, but different from that of $S^{2}\times S^{2}$ and the Page metric, both of which exhibit one negative mode Volkov:2000ih ; Hennigar:2024gbg .
@@
$\tilde{\mathcal{I}}_{21}=\int_{\varepsilon}^{\infty}\frac{dt}{\sqrt{t^{2}-\varepsilon^{2}}}\frac{-8e^{-2t}+31e^{-t}+27e^{t}-44}{\left(e^{t}-1\right)^{4}}\times\cosh\left(2\sqrt{t^{2}-\varepsilon^{2}}\right)~.$ (6.32)
@@
$\tilde{\mathcal{B}}_{21}=-\frac{599}{120}~.$ (6.33)
The heat kernel integral $\tilde{\mathcal{I}}_{21}$ is evaluated in appendix B.2, and the resulting expression is provided in (B.4).
@@
$\tilde{\mathcal{I}}_{22}=\int_{\varepsilon}^{\infty}\frac{dt}{\sqrt{t^{2}-\varepsilon^{2}}}\frac{-20e^{-\frac{5t}{2}}+80e^{-\frac{3t}{2}}-118e^{-\frac{t}{2}}+70e^{\frac{t}{2}}}{\left(e^{t}-1\right)^{4}}\times\cosh\left(\tfrac{\sqrt{13}}{2}\sqrt{t^{2}-\varepsilon^{2}}\right)~.$ (6.34)
@@
$\tilde{\mathcal{B}}_{22}=-\frac{607}{30}~.$ (6.35)
The heat kernel integral $\tilde{\mathcal{I}}_{22}$ is evaluated in appendix B.2, and the resulting expression is provided in (B.6).
@@
$\tilde{\mathcal{I}}_{23}=\int_{\varepsilon}^{\infty}\frac{dt}{\sqrt{t^{2}-\varepsilon^{2}}}\frac{4e^{-t}\left(9\sinh t+19\cosh t-16\right)}{\left(e^{t}-1\right)^{4}}\times\cosh\left(\sqrt{t^{2}-\varepsilon^{2}}\right)~.$ (6.36)
@@
$\tilde{\mathcal{B}}_{23}=\frac{1741}{60}~.$ (6.37)
The heat kernel integral $\tilde{\mathcal{I}}_{23}$ is evaluated in appendix B.2, and the resulting expression is provided in (B.8).
@@ 6.5 Comparing to the heat kernel coefficient
We note that, as is indicated in the $\Lambda_{\text{u.v.}}$ dependent pre-factor of (6.21), the final $2\times 8+1$ stems from a proper treatment of the path integration measure. A similar treatment can be done for the linearised Einstein theory about the round $S^{4}$ saddle, see for example Volkov:2000ih ; Anninos:2020hfj , which yields a coefficient
@@
where now $\text{vol}_{\text{can}}=\tfrac{2}{3}(2\pi)^{6}$ is the canonically normalised group volume of the $SO(5)$ isometry group of $S^{4}$. The overall phase of $\mathcal{Z}^{(1)}_{S^{4}}$ is $(\pm i)^{1+5}=-1$ Polchinski:1988ua .
@@ Acknowledgements
It is a great pleasure to acknowledge discussions with Alejandra Castro, Damián Galante, Pietro Benetti Genolini, Thomas Hertog, Chris Herzog, Albert Law, Juan Maldacena, Miguel Montero, Beatrix Mühlmann, and Edgar Shaghoulian. We are very grateful to Themistoklis Zikopoulos for pointing out the volume of $V_{\mathbb{C}P^{2}\#2\overline{\mathbb{C}P^{2}}}$. D.A. is funded by the Royal Society under the grant “Concrete Calculables in Quantum de Sitter” and the STFC Consolidated grant ST/X000753/1. C.B. is funded by STFC under the grant reference STFC/2887726. S.B. is funded by STFC under the grant reference STFC/2928477.
@@ A.1 Comparison to coincident point limit of Green’s function
$\frac{1}{2}\times V_{\mathbb{C}P^{2}}\times\lim_{\sigma\to 1}\partial_{\mu^{2}}^{(3)}G_{\mathbb{C}P^{2}}(\sigma)|_{\mu^{2}=1}=\partial_{\mu^{2}}^{(4)}\log\mathcal{Z}^{(\varphi)}_{\mathbb{C}P^{2}}|_{\mu^{2}=1}~.$ (A.9)
@@ A.2 Integral representation of Green’s function
In the above, X{X} and Y{Y} are unit normal vectors on $\mathbb{R}^{6}$, $\sigma\equiv{X}\cdot{Y}$, and now we have
@@
$\frac{1}{2}\times V_{\mathbb{C}P^{2}}\times\lim_{\sigma\to 1}\partial_{\mu^{2}}^{(3)}G_{\mathbb{C}P^{2}}(\sigma)|_{\mu^{2}=4}=\frac{3\zeta(5)}{256}~,$ (A.15)
@@
It is interesting that the expression (A.14) takes a character-like form resembling the expressions in Anninos:2020hfj for the sphere. To do so it is convenient to use that $\lambda\partial_{\lambda}\partial_{\nu}\tfrac{\lambda^{\pm i\nu}}{\log\lambda}=\nu\lambda^{\pm i\nu}$. Denoting ${\boldsymbol{\Delta}}=2+2i\nu$, with $\nu=\sqrt{\tfrac{1}{4}\mu^{2}-1}$, and upon integrating by parts in $\lambda$, we can re-express (A.14) as
$\frac{1}{2}\times V_{\mathbb{C}P^{2}}\times G_{\mathbb{C}P^{2}}(\zeta^{i},\zeta^{i})=\frac{1}{8\nu}\partial_{\nu}\int_{0}^{1}\frac{d\lambda}{2\lambda(-\log\lambda)}\frac{1+4\lambda+\lambda^{2}}{1-\lambda}\frac{\lambda^{\frac{\boldsymbol{\Delta}}{2}}+\lambda^{\frac{\bar{\boldsymbol{\Delta}}}{2}}}{(1-\lambda)^{3}}~.$ (A.16)